OpenStudy (anonymous):

Fans & Medals!! Complete the equation for the linear function whose graph contains the points (9, 7) and (4, –8). y −____=____(x − 9)

OpenStudy (anonymous):

Find the slope first m=(y2-y1)/(x-x1) and then plug in points x1 and y1...your equation is y-y1=m(x-x1)

OpenStudy (anonymous):

I dunno. im only im 5th grade

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

I got it -15/-5 @iambatman